# Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction

### v4

@@ Abstract
Commodity futures are volatile. Forecasting across horizons with interpretable drivers remains challenging. We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction latent-variable model that combines sparse coding, unrolled optimization, and amortized inference. SLFF explicitly optimizes a sparse latent code to explain multi-horizon futures trajectories and trains an encoder whose outputs are validated against the optimization-based solution before deployment. The method is paired with an information-set-aware pipeline (vintage macro releases, lag-aware fills, leakage checks) and evaluated under rolling-origin folds against representative statistical and neural baselines. We provide quantitative criteria for factor labeling and directional diagnostics that account for no-change regimes. On Copper and WTI futures (2005–2023), SLFF achieves competitive RMSE/MAE, improves directional skill beyond persistence, and yields factors that are stable across seeds and linked to measurable fundamentals. Code, diagnostics, and information-set specifications are released for reproducibility.
TITLE PAGE
Manuscript Title:
Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction
Manuscript Type:
Research Article
Author:
Abhijit Gupta, PhD
Affiliation: Independent Researcher
ORCID: 0000-0002-6292-3789
Email: abhijit.gupta_in@outlook.com
Corresponding Author:
Abhijit Gupta, PhD (email above)
Running Title (suggested):
SLFF for Commodity Futures Forecasting
Keywords (suggested):
commodity forecasting; multi-horizon prediction; latent factors; interpretability; time series; deep learning
Word Count (approx., excluding references):
7204
Figures:
4
Tables:
16
Funding:
None
Competing Interests:
None
Ethics Approval:
Not applicable (no human or animal subjects).
Data Availability:
See the manuscript Data Availability and Licensing section. The study uses a mix of public and proprietary sources; proprietary datasets require the corresponding licenses.
Permissions to Reproduce Material:
All figures and tables are original. Underlying data include public and proprietary sources; permissions/licenses are as described in the Data Availability section.
@@ 1 Introduction
Commodity markets, and their associated futures contracts, form a critical part of the global economy, providing essential raw materials and hedging instruments. The prices of key commodity futures, such as those for industrial metals like Copper and energy benchmarks like Crude Oil, are not merely transactional figures; they are vital indicators of economic health, drivers of industrial costs, and significant factors influencing inflation, trade balances, and geopolitical dynamics. Copper futures, for instance, are closely watched for signals about global industrial demand. Crude Oil futures are pivotal for managing energy price risk and reflect complex geopolitical and supply/demand interactions.
The inherent volatility of these commodity futures prices presents substantial risks and opportunities for a wide spectrum of market participants. Producers and consumers use futures to hedge price uncertainty, while financial institutions and traders engage in speculative and arbitrage activities. Consequently, the ability to accurately forecast commodity futures prices across various time horizons is of paramount importance for effective risk mitigation, optimized resource allocation, and informed investment decisions.
The challenge lies in the sheer complexity of factors driving futures prices. These markets are influenced by a confluence of interconnected forces:
@@
    Macroeconomic Environment: Global GDP growth, industrial production, interest rates, currency exchange rates (particularly the US Dollar), and inflation expectations.
@@
    Underlying Physical Market Fundamentals: For Copper, this includes mine production, scrap availability, and inventory levels. For Crude Oil, factors include OPEC+ decisions, non-OPEC output, geopolitical events in producing regions, and inventory levels.
@@
    Demand Fundamentals: Linked to manufacturing activity, construction, transportation needs, and energy transition dynamics.
  4. 4.
    Market Sentiment and Financialization: Speculative positioning in futures and options markets (e.g., CFTC Commitment of Traders reports), algorithmic trading, and broader risk-on/risk-off sentiment.
  5. 5.
    Other factors: Including weather patterns impacting supply/demand and inter-commodity relationships.
This intricate web of high-dimensional, often collinear, and dynamically interacting drivers makes futures price forecasting an exceptionally challenging task, characterized by non-linear relationships and regime shifts. Existing approaches often face a trade-off between predictive accuracy and interpretability. Classical time series models (ARIMA, GARCH) may struggle with non-linearities and exogenous factors. Econometric structural models rely on strong assumptions. While traditional machine learning (SVM, Random Forests) and standard deep learning models (LSTMs, CNNs) can capture complex patterns, they often function as ”black boxes,” limiting actionable insights [41]. There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
Given these limitations, this research addresses the question: How can we develop a latent-variable forecasting framework that remains faithful to the information set, attains robust multi-horizon accuracy, and exposes stable economic factors? We answer with the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction model that uses an explicit latent-factor optimization layer together with a LISTA-style amortized encoder [34]. The latent code $\mathbf{z}$ is never treated as a reconstruction target; it is the solution of a constrained regression problem that explains future trajectories under an L1 sparsity prior.
This paper makes the following contributions:
  1. 1.
    Method: We introduce SLFF, a sparse latent-factor forecaster with iterative inference, and describe a latent inference objective that supports both test-time amortization and (optionally) unrolled, differentiable optimization.
  2. 2.
    Information set: We specify an information-set-aware data pipeline, including release calendars and vintage macroeconomic series (ALFRED/FRED-MD, IMF WEO vintages), and provide an “available-at-tt” specification to prevent mixed-frequency leakage.
  3. 3.
    Evaluation: We evaluate on six rolling-origin folds (train: 6y, val: 1y, test: 1y) with test years spanning 2013–2023 (every two years), reporting mean $\pm$ standard deviation across folds and random seeds.
  4. 4.
    Baselines: We compare against PatchTST, Temporal Fusion Transformer (TFT), N-HiTS, DeepAR, and tuned gradient boosting models under matched hyperparameter budgets and learning-rate schedules.
  5. 5.
    Ablations: We present ablations over sparsity, iterative inference, latent dimensionality, and training regimes to quantify their impact on deployed forecasting accuracy and latent interpretability.
  6. 6.
    Directional diagnostics: We report directional metrics that account for no-change regimes (no-change rate, class-conditional hit rates, MCC, calibration curves, confusion matrices).
  7. 7.
    Factor validation: We evaluate factor stability via Procrustes alignment across random seeds and validate factors against observable disruption/event indicators, including counterfactual perturbation tests.
  8. 8.
    Reproducibility: We provide code (JAX/Flax), experiment manifests, Dockerized data pipelines, compute footprints, and an executable feature notebook to support replication and extension.
The remainder of this paper revisits the related work, describes SLFF in detail, specifies the curated information set, explains the experimental protocol, and presents empirical/interpretability evidence before concluding with limitations and future work.
@@ 2 Review of Related Work in Commodity Forecasting
2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods affect economic activities and individual well-being [30]. Macroeconomic stability of commodity-dependent countries is closely linked to price fluctuations [15], and predicting staple commodity prices is vital for food security [29]. Reliable forecasts are essential for producers, traders, investors, and policymakers.
Commodity price forecasting is notably challenging due to inherent market complexity, volatility, nonlinear relationships, and nonstationary trends [37]. Prices respond to numerous interrelated factors including supply/demand fundamentals, macroeconomic indicators, financial market influences, geopolitical events, and environmental variables. These factors evolve, with increased financialization, global crises, and climate change adding unpredictability, often rendering traditional methods inadequate.
2.2 Traditional Econometric Approaches: Foundations and Limitations
Quantitative commodity forecasting historically relied on models like ARIMA, VAR, and VECM. Box-Jenkins ARIMA models [11] gained popularity for their statistical properties [64]. GARCH models [10, 27] were developed to address time-varying volatility. However, these models struggle with the nonlinearities and structural breaks common in commodity prices [84, 73], often underperforming [9]. Their limitations spurred exploration of more flexible paradigms [75, 52].
2.3 Rise of Machine Learning in Commodity Forecasting
ML and AI techniques have become prominent alternatives [16, 80]. ML models inherently handle complex, nonlinear relationships without restrictive assumptions, often improving accuracy over traditional methods [65]. They can also integrate diverse data, including unstructured text via NLP [91]. This data-driven approach [13] contrasts with theory-driven econometrics but typically requires large datasets, a potential constraint in some markets [32].
2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied (Table 1).
Table 1: Overview of Machine Learning Models in Commodity Forecasting.
Model Type Core Principle Strengths in Commodity Forecasting Weaknesses / Challenges
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Prone to overfitting if not properly tuned, less interpretable than single trees.
Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons. Universal approximators, capture non-linear patterns. Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
Long Short-Term Memory (LSTM) Type of RNN with specialized gating mechanisms (input, forget, output gates) to control information flow and capture long-range dependencies in sequential data. Excels at modeling sequential data and long-term dependencies, mitigating vanishing gradient problem of simple RNNs. Computationally expensive, requires significant data, complex architecture demanding careful tuning.
Gated Recurrent Unit (GRU) Simplified LSTM version with fewer gates (reset and update gates), also for long-range dependencies. Captures long-term dependencies like LSTM, potentially more computationally efficient due to fewer parameters. Performance relative to LSTM is context-dependent, may not always match LSTM’s capabilities on complex tasks.
Transformer Uses self-attention mechanisms to weigh input sequence importance globally, allowing parallel processing of sequence elements. Excellent at capturing very long-range dependencies, often outperforming RNNs/LSTMs on such tasks; parallelizable. Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
Note: The table provides a compact overview; details and examples are given in the cited references.
Foundational models include SVR [25], noted for handling non-linearities [87, 47, 71, 70], and tree-based ensembles like RF [12] and XGBoost [18], valued for robustness and interaction modeling [86].
Deep learning (DL) advanced from ANNs/MLPs [79, 93] to RNNs [26]. LSTMs [42] and GRUs [19] became key for sequential data due to their ability to learn long-term dependencies by addressing the vanishing gradient issue [8]. They are widely used in commodity forecasting [28, 95, 83, 92, 20], with variants like bidirectional LSTMs [82] and CNN-LSTM hybrids [50, 61, 54, 7]. The choice involves balancing complexity, data needs, and interpretability [40].
Transformer models [88], leveraging self-attention, represent a significant advancement for long-range dependency capture [58, 57, 60]. Recent architectures like PatchTST [68] and FEDformer [96] highlight ongoing progress. Transformers are often integrated into hybrid models, frequently with decomposition methods [89]. Large Language Models (LLMs) are also influencing financial forecasting research [14, 90].
Hybrid models are prevalent, often combining traditional methods with ML, or using a ”decomposition and ensemble” strategy [94, 49, 53]. Decomposition techniques (e.g., Wavelets [22], VMD [24], EMD [43], SSA [31], STL [21]) simplify series for component-wise forecasting [6, 35, 39, 56]. Integration of XAI is also an emerging trend [72].
Performance is typically assessed using RMSE, MAE, and directional accuracy [45]. Statistical tests like Diebold-Mariano [23] are sometimes used for comparing model accuracy. While ML/DL models often show superior performance, rankings are context-dependent [85, 46].
@@ 3 Methodology: Sparse Latent Factor Forecaster with Iterative Inference
This section details the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from structured prediction energy networks [55], sparse coding [69], and differentiable optimization layers [3], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
@@ 3.3 Latent Energy and Iterative Inference
We define
$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2},$ (1)
$C(Y_{t},\tilde{Y}_{t})$ $=\frac{1}{N}\|Y_{t}-\tilde{Y}_{t}\|_{2}^{2},$ (2)
where $\lambda$ enforces sparsity and $\mu$ ties inference to the amortized encoder. We denote the optimization-refined latent by
$\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z}),$ (3)
which is a constrained refinement of $\text{Enc}_{X\rightarrow Z}(X_{t})$. For clarity, the unconstrained oracle (not used in training or evaluation) corresponds to the special case $\mu=0$:
$\mathbf{z}^{\dagger}_{t}=\arg\min_{\mathbf{z}}\;C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}.$ (4)
We solve the refinement with $K$ steps of proximal gradient descent,
$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha(\nabla_{\mathbf{z}}C(Y_{t},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}_{t}))+2\mu(\mathbf{z}^{(k)}-\text{Enc}_{X\rightarrow Z}(X_{t})))\big),$ (5)
where the shrinkage operator is applied element-wise. In practice we warm-start inference at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to encourage predictable latent codes; unless otherwise stated, this includes the $\mu=0$ setting, so the $\mu=0$ sweep row reflects a warm-started refinement rather than a cold-start oracle. In the $\mu$-sweep we tune $\alpha$ per $\mu$ to ensure stable convergence (monotone energy decrease), so the sweep is not confounded by step-size instability. The iterations are differentiable, enabling end-to-end training or truncated backpropagation. When enabled (not in the main experiments), the contrastive regularizer described in Section 5 is applied only during encoder training and does not modify the inference energy.

### v5

@@ Abstract
Amortized variational inference in latent-variable forecasters creates a deployment gap: the test-time encoder approximates a training-time optimization-refined latent, but without access to future targets. This gap introduces unnecessary forecast error and interpretability challenges. In this work, we propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), addressing this through (i) a sparse coding objective with L1 regularization for low-dimensional latents, (ii) unrolled proximal gradient descent (LISTA-style) for iterative refinement during training, and (iii) encoder alignment to ensure amortized outputs match optimization-refined solutions. Under a linearized decoder assumption, we derive a design-motivating bound on the amortization gap based on encoder-optimizer distance, with convergence rates under mild conditions; empirical checks confirm the bound is predictive for the deployed MLP decoder. To prevent mixed-frequency data leakage, we introduce an information-set-aware protocol using release calendars and vintage macroeconomic data. Interpretability is formalized via a three-stage protocol: stability (Procrustes alignment across seeds), driver validity (held-out regressions against observables), and behavioral consistency (counterfactuals and event studies). Using commodity futures (Copper, WTI, Gold; 2005–2025) as a testbed, SLFF demonstrates significant improvements over neural baselines at 1- and 5-day horizons, yielding sparse factors that are stable across seeds and correlated with observable economic fundamentals (interpretability remains correlational, not causal). Code, manifests, diagnostics, and artifacts are released.
@@ 1 Introduction
Latent-variable forecasters decompose a high-dimensional time series into a lower-dimensional latent representation that summarizes historical context and predicts future values. The standard approach—amortized variational inference—trains an encoder network to map observations to latent codes such that a decoder can predict targets. However, during training, access to targets $Y$ enables test-time-inaccessible optimization of the latent code; at deployment, only the encoder output is available. This mismatch between training-time iterative refinement and test-time amortized inference creates the amortization gap [2, 13]. The gap degrades both forecast accuracy and interpretability, because (i) the test-time encoder may not learn to emulate the iterative solution, and (ii) sparse latent factors discovered under iterative refinement may become dense or unstable when collapsed to amortized form.
Three prior approaches attempt to bridge this gap: (i) Variational Autoencoders (VAEs) enforce a prior on the latent space, but the variational approximation and KL regularization are optimized for reconstruction, not forecasting; (ii) Semi-amortized variational inference [8] uses fixed-point iterations on the encoder mean at test time, but this adds latency incompatible with real-time forecasting; (iii) Test-time optimization fine-tunes latents using the first few observed target steps, but requires causal structure knowledge and breaks when labels are unavailable. For forecasting specifically, the challenge is acute: the latent must explain not a reconstruction $\hat{X}$ but a predictive target $Y$ that is not observable at test time. Existing methods do not explicitly regulate the encoder-to-optimizer distance, nor do they provide theoretical guarantees on the incurred gap.
We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), which directly minimizes the amortization gap through three coupled design choices. First, the latent is defined not as a generative variable but as a structured prediction variable—the explicit solution to a sparse regression problem that explains the target trajectory under an L1 sparsity prior. Second, iterative inference is not an optional post-hoc refinement; it is an integral part of the training loop. We unroll proximal gradient descent (inspired by LISTA) and backpropagate through the refinement steps, allowing gradients to flow from the forecasting loss into the optimization process. Third, the amortized encoder is trained not to match an arbitrary latent prior but to match the optimization-refined solution. During training, we compute the refined latent $\mathbf{z}^{\star}$ conditioned on the true target $Y$, then minimize the encoder’s squared distance to $\mathbf{z}^{\star}$ in a two-stage loop. Crucially, we include a proximity term in the inference objective ($\mu\|\mathbf{z}-\text{Enc}(X)\|_{2}^{2}$) that keeps the optimization path predictable from $X$ alone, ensuring the training-time refinement is a plausible target for the encoder. The method provides an amortization gap bound showing how the test-time forecast error depends on the encoder-optimizer distance.
Commodity futures markets offer a natural testbed for latent-variable forecasting methods. They exhibit high-dimensional input spaces (macroeconomic indicators, supply/demand fundamentals, financial flows, technical signals), mixed-frequency releases (daily exchange prices, weekly inventories, monthly PMIs, quarterly GDP), and economically important predictions (hedging, risk management, portfolio construction). Unlike low-frequency macroeconomic series, commodity prices move daily with observable signals, enabling rapid validation of both forecast accuracy and factor interpretability. Moreover, commodity markets are efficiently priced but not fully predictable, sitting at the frontier of forecasting difficulty: classical models and persistence baselines are strong, but commodity-specific structural models (supply curves, convenience yields) provide weak competition when translated to the machine-learning setting.
This paper makes three core contributions:
@@
    Methodological contribution: SLFF with theoretical analysis. We introduce SLFF as a latent-variable forecaster that explicitly optimizes a sparse code under L1 regularization, trains an encoder to match the optimization-refined solution, and provides (i) an upper bound on the amortization gap in terms of encoder-optimizer alignment, (ii) convergence rates for the proximal gradient iterations, and (iii) conditions under which the deployed (amortized) path achieves near-oracle performance. This bridges the theoretical gap between training-time iterative inference and test-time amortized prediction.
@@
    Information-set protocol: Vintage-aware data alignment. We formalize an information-set-aware evaluation protocol that prevents mixed-frequency data leakage. We use release calendars, ALFRED-style vintage repositories, and lag-aware forward-filling to ensure that the feature matrix $X_{t}$ contains only information available at end-of-day $t$. This protocol is absent from most deep-learning forecasting papers and is essential for avoiding false discoveries in commodity markets.
@@
    Interpretability protocol: Three-stage validation. We introduce a structured protocol for validating sparse latent factors: (i) stability, measured by Procrustes-aligned subspace overlap across random seeds and folds; (ii) driver validity, measured by regression on held-out observable indices not in the decoder’s feature set; (iii) behavioral consistency, measured by counterfactual perturbations and event-window analysis. This protocol moves beyond attention visualization and SHAP values toward systematic, quantitative factor validation.
The remainder of the paper proceeds as follows. Section 2 reviews latent-variable models in forecasting, sparse coding and dictionary learning, amortized versus optimization-based inference, and interpretability in financial machine learning. Section 3 presents the SLFF architecture, energy formulation, iterative refinement procedure, and training algorithm in detail. Section 5 provides theoretical analysis with propositions bounding the amortization gap and convergence rates. Section 4 specifies the information set, data preprocessing, and rolling-origin evaluation protocol. Section 6 reports predictive accuracy, alignment diagnostics, ablations, and interpretability analysis across Copper, WTI, and Gold futures. Section 7 provides a framework for understanding SLFF’s applicability beyond the three test commodities. Section 8 discusses limitations and future directions. Code, reproducibility artifacts, and diagnostic notebooks are released.
@@ 2 Review of Related Work in Commodity Forecasting
2.1 Latent-Variable Models in Forecasting
Latent-variable approaches to time-series forecasting aim to discover low-dimensional representations that explain complex dynamics. Classical state-space models [6] use unobserved components (trend, seasonality, latent factors) combined with linear dynamics and observation equations. More recent work extends this with neural encoders: deep state-space models [16] replace linear dynamics with RNNs; VAE-based forecasters [18] add an amortized encoder and latent prior; structured inference networks [9] combine amortized and iterative inference. SLFF builds on the structured inference idea but specialized for forecasting: the latent is a sparse code explained by a target, not a reconstruction target, and the encoder is trained to match the optimization-refined code rather than approximate an arbitrary prior.
2.2 Sparse Coding and Dictionary Learning
Sparse coding [15] models data as a sparse linear combination of basis vectors (a dictionary), with solutions computed via L1-regularized optimization. Mairal et al. [12] extend sparse coding to online/streaming settings. In forecasting, sparsity improves interpretability: a forecast explained by a few latent factors is more actionable than one driven by dozens of dense hidden units. LISTA (Learned Iterative Soft-Thresholding Algorithm) [5] unrolls proximal gradient descent as a neural network, enabling end-to-end learning of both the dictionary and the inference procedure. SLFF adopts LISTA’s unrolled structure but applies it within a forecasting objective, with explicit encoder alignment to ensure the learned iterative solution generalizes.
2.3 Amortized versus Optimization-Based Inference
Amortized inference maps observations to latent codes via a fixed encoder, achieving O(1) test-time cost but potentially sacrificing accuracy. Optimization-based inference refines latents iteratively given data and targets, with higher accuracy but O(K) cost for $K$ steps. The tension is formalized in the amortization gap literature [2, 13]: the test-time encoder cannot match the train-time iterative solution if the latter depends on information unavailable at deployment. Semi-amortized variational inference [8] uses fixed-point iteration at test time; structured prediction energy networks [1] add optimization layers to deep learning. SLFF addresses the gap directly by (i) keeping test-time refinement tractable via a proximity term to the encoder, (ii) providing a theoretical bound on the gap, and (iii) validating that the deployed amortized path achieves competitive accuracy.
2.4 Interpretability in Forecasting
Attention-based forecasters (e.g., Temporal Fusion Transformer [10]) expose input importance through attention weights. SHAP [11] provides local feature attribution. Traditional factor models (PCA, ICA, NMF) extract latent structure but often lack predictive power; recent work [17] focuses on economic interpretability. SLFF’s sparse latent factors inherit benefits of classical factors (low-dimensional, interpretable) but are learned end-to-end for forecasting. The three-stage interpretability protocol (stability, driver validity, behavioral consistency) provides quantitative validation missing from attention-visualization-only approaches.
@@ 3 Methodology: Sparse Latent Factor Forecaster with Iterative Inference
Notation
Table 1: Key notation used in the methodology section.
Symbol Definition
$\mathbf{x}_{t}$ Feature vector available at end of day $t$
$X_{t}$ Input tensor of look-back window ($W\times d$)
$Y_{t}$ Target future log prices at horizons $\tau_{1},\ldots,\tau_{N}$
$\mathbf{z}_{t}$ Latent code (sparse, $\mathbb{R}^{m}$)
$\mathbf{h}_{t}$ History context from $\text{Pred}(X_{t})$
$\text{Enc}_{X\to Z}$ Amortized encoder network
Dec Decoder mapping latent and context to forecasts
$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ Energy function for optimization
$\mathbf{z}_{t}^{\star}$ Optimization-refined latent code
@@ 3.3 Latent Energy and Iterative Inference
The training objective combines predictive loss and matching:
$L=L_{\text{pred}}+\beta L_{\text{match}},$ (1)
where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$ with $C$ as mean squared error over horizons, and $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$. The energy function is:
$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}.$ (2)
Optimization yields $\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z})$. We solve via proximal gradient descent with $K=10$ iterations:
$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha\nabla_{\mathbf{z}}[C+\mu\|\mathbf{z}^{(k)}-\text{Enc}\|_{2}^{2}]\big),$ (3)
where SoftThresh applies element-wise shrinkage, $\alpha=0.01$ is the step size, and initialization is warm-started at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to promote consistency. Gradients are computed via automatic differentiation, enabling end-to-end training.